# Mini-Game Lifetime Value Prediction in WeChat

### v2

@@ 1. Introduction
To address the aforementioned challenges, we propose Graph-Represented Pareto-Optimal LifeTime Value prediction (GRePO-LTV). Our method adopts graph representation learning to alleviate the paucity of purchase records and build effective user and game embeddings. By learning graph-based representations that capture the complex interactions between users and games, we can better encode the collaborative signal (Wang et al., 2019) and utilize the limited available data to improve LTV prediction accuracy. Furthermore, our method employs Pareto optimization to avoid task contradiction between value predictions of different time horizons. This multi-objective optimization approach allows us to find a set of Pareto-optimal solutions that balance the trade-offs between short-term and long-term prediction accuracy. By considering the Pareto front, we can select a model that achieves strong performance across all time horizons without compromising on either end of the spectrum.
@@ 2.1. Problem Definition
In formal terms, there exists a set of users $\mathcal{U}=\{u_{p}\}_{p=1}^{P}$, accompanied by a corresponding feature set $\mathcal{X}_{u}=\{x_{u_{p}}\}_{p=1}^{P}$ and a sequence of behavioral patterns $\mathcal{B}_{u}=\{b_{u_{p}}\}_{p=1}^{P}$. Additionally, there is a set of items =$\mathcal{I}=\{i_{q}\}_{q=1}^{Q}$, endowed with its own feature set $\mathcal{X}_{i}=\{x_{i_{q}}\}_{q=1}^{Q}$. Each $b_{u_{p}}$ is constituted by an array of items that have previously been interacted with by the user $u_{p}$.
@@ 3.1. Graph Representation Learning
where $\mathcal{V}_{\mathrm{masked}}$ represents the set of masked nodes, $x_{v}$ and $\hat{x}_{v}$ denote the attribute vectors of node $v$ and their reconstructed versions after masking, respectively, and $\xi_{a}$ is a scaling coefficient for the attribute reconstruction loss. We combine these two losses using a hyperparameter $\zeta$:
@@ 3.2. Backbone of GRePO-LTV
The tower layer integrates and processes various embeddings to generate final predictions. It employs AdaSparse (Yang et al., 2022) for domain-specific neuron filtering, adapting the network structure for each domain. To handle the long-tailed distribution of user values, we adopt a zero-inflated lognormal distribution (Wang et al., 2019) modeling approach. We refer our readers to Appendix B for detailed explanations about ZILN.
@@ 3.3. Pareto Optimization
[…]
@@
$\left\{\begin{array}[]{lr}\lambda_{1}=\sin{\phi}\cos{\theta},&\\ \lambda_{2}=\sin{\phi}\sin{\theta},\\ \lambda_{3}=\cos{\phi}.&\end{array}\right.$
@@ 4.3. Baselines
⑤ ZILN (Wang et al., 2019) is a zero-inflated lognormal loss function to address the imbalanced problem for LTV prediction.
@@ B.1. Overview
The Zero-Inflated Lognormal (ZILN) distribution is a statistical model designed for datasets that exhibit a high frequency of zero values alongside positive, continuous data that are typically right-skewed (Wang et al., 2019). Standard continuous probability distributions, including the conventional lognormal distribution, are often inadequate for such data as they cannot inherently model a significant point mass at zero (Crow and Shimizu, 1987). The ZILN model addresses this by conceptualizing the data as arising from a mixture of two distinct processes: one that generates zero values and another that generates positive values following a lognormal distribution.
@@ B.4. Mathematical Definition
Let $Y$ be a random variable following a Zero-Inflated Lognormal distribution. The distribution is characterized by three parameters: the zero-inflation probability $\pi$ ($10\leq\pi<$ 1), and the lognormal parameters $\mu$ (mean of $\ln Y$ for $Y>0$) and $\sigma^{2}$ (variance of $\ln Y$ for $Y>0$, with $\sigma>0$). The probability density function (PDF) of the ZILN distribution, $g(y;\pi,\mu,\sigma^{2})$, is expressed in a piecewise form as:

### v3

@@ 1. Introduction
To address the aforementioned challenges, we propose Graph-Represented Pareto-Optimal LifeTime Value prediction (GRePO-LTV). Our method adopts graph representation learning to alleviate the paucity of purchase records and build effective user and game embeddings. By learning graph-based representations that capture the complex interactions between users and games, we can better encode the collaborative signal (Wang et al., 2019a) and utilize the limited available data to improve LTV prediction accuracy. Furthermore, our method employs Pareto optimization to avoid task contradiction between value predictions of different time horizons. This multi-objective optimization approach allows us to find a set of Pareto-optimal solutions that balance the trade-offs between short-term and long-term prediction accuracy. By considering the Pareto front, we can select a model that achieves strong performance across all time horizons without compromising on either end of the spectrum.
@@ 2.1. Problem Definition
In formal terms, there exists a set of users $\mathcal{U}=\{u_{p}\}_{p=1}^{P}$, accompanied by a corresponding feature set $\mathcal{X}_{u}=\{x_{u_{p}}\}_{p=1}^{P}$ and a sequence of behavioral patterns $\mathcal{B}_{u}=\{b_{u_{p}}\}_{p=1}^{P}$. Additionally, there is a set of items $\mathcal{I}=\{i_{q}\}_{q=1}^{Q}$, endowed with its own feature set $\mathcal{X}_{i}=\{x_{i_{q}}\}_{q=1}^{Q}$. Each $b_{u_{p}}$ is constituted by an array of items that have previously been interacted with by the user $u_{p}$.
@@ 3.1. Graph Representation Learning
where $\mathcal{V}_{\mathrm{msk}}$ represents the set of masked nodes, $x_{v}$ and $\hat{x}_{v}$ denote the attribute vectors of node $v$ and their reconstructed versions after masking, respectively, and $\xi_{a}$ is a scaling coefficient for the attribute reconstruction loss. We combine these two losses using a hyperparameter $\zeta$:
@@ 3.2. Backbone of GRePO-LTV
The tower layer integrates and processes various embeddings to generate final predictions. It employs AdaSparse (Yang et al., 2022) for domain-specific neuron filtering, adapting the network structure for each domain. To handle the long-tailed distribution of user values, we adopt a zero-inflated lognormal distribution (Wang et al., 2019b) modeling approach. We refer our readers to Appendix B for detailed explanations about ZILN.
@@ 3.3. Pareto Optimization
$\text{ s.t. }\boldsymbol{\beta}^{\top}G^{\top}g_{j}\geqslant 0\quad\forall j\in\mathrm{J}=\left\{\begin{array}[]{cc}\mathrm{J}^{*}&\mathrm{KL}\left(\mathcal{L}_{\mathrm{LTV}}\odot\boldsymbol{\lambda}|\boldsymbol{1}\right)\leqslant\epsilon\\ {[m]}&\mathrm{KL}\left(\mathcal{L}_{\mathrm{LTV}}\odot\boldsymbol{\lambda}|\boldsymbol{1}\right)>\epsilon\end{array},\right$.
$\text{ where }\quad\mathrm{J}^{*}=\left\{j\in[m]\mid j=\arg\max_{j^{\prime}\in[m]}l_{j^{\prime}}\lambda_{j^{\prime}}\right\},$
@@
[…]
@@ 4.3. Baselines
⑤ ZILN (Wang et al., 2019b) is a zero-inflated lognormal loss function to address the imbalanced problem for LTV prediction.
@@ B.1. Overview
The Zero-Inflated Lognormal (ZILN) distribution is a statistical model designed for datasets that exhibit a high frequency of zero values alongside positive, continuous data that are typically right-skewed (Wang et al., 2019b)f. Standard continuous probability distributions, including the conventional lognormal distribution, are often inadequate for such data as they cannot inherently model a significant point mass at zero (Crow and Shimizu, 1987). The ZILN model addresses this by conceptualizing the data as arising from a mixture of two distinct processes: one that generates zero values and another that generates positive values following a lognormal distribution.
@@ B.4. Mathematical Definition
Let $Y$ be a random variable following a Zero-Inflated Lognormal distribution. The distribution is characterized by three parameters: the zero-inflation probability $\pi$ ($0\leq\pi<1$), and the lognormal parameters $\mu$ (mean of $\ln Y$ for $Y>0$) and $\sigma^{2}$ (variance of $\ln Y$ for $Y>0$, with $\sigma>0$). The probability density function (PDF) of the ZILN distribution, $g(y;\pi,\mu,\sigma^{2})$, is expressed in a piecewise form as: